OpenStudy (yacoub1993):

Use the formula for the sum of the first n terms of a geometric sequence to solve. Find the sum of the first 8 terms of the geometric sequence: -8, -16, -32, -64, -128, . . . .

hartnn (hartnn):

know the formula ?

OpenStudy (yacoub1993):

no i dont know

OpenStudy (yacoub1993):

@hartnn Help me solve it

hartnn (hartnn):

\(\large S_n = a_1 \dfrac{r^n -1 }{r-1}\)

hartnn (hartnn):

where a1 i s the first term = ...? n = number of terms = 8 is given r = common ratio = ....? can u find ?

OpenStudy (yacoub1993):

ok what is the common ratio

hartnn (hartnn):

its the ratio of next term/ current term r = 2nd term/1st term = 3rd term/2nd term = 4th term/3d term =..... and so on

hartnn (hartnn):

so whats r here ?

OpenStudy (yacoub1993):

r is 8 right

hartnn (hartnn):

no.... r = -16/-8 = 16/8 = ... ?

OpenStudy (yacoub1993):

2

OpenStudy (yacoub1993):

when i add them up all four after dividing =8

hartnn (hartnn):

yes, r=2 a1 = 1st term =-8 n=8 just plug these in the formula!

hartnn (hartnn):

why did u add them up ?

OpenStudy (yacoub1993):

if i dont add them up they are all equal 2

hartnn (hartnn):

did u get how r =2 now ?

OpenStudy (yacoub1993):

yes

hartnn (hartnn):

good, so use the formula, u have all the values now

OpenStudy (the_fizicx99):

you can pretty much just tell what the ratio is by looking at the sequence :)

OpenStudy (yacoub1993):

i got as final answer \[S _{8}=-8\frac{ 2^{8}-1 }{ 2-1 }\] \[S _{8}=-1024\]

OpenStudy (yacoub1993):

how can i further simplify it

hartnn (hartnn):

whats 2^8 =... ?

OpenStudy (yacoub1993):

it was r^n right

hartnn (hartnn):

yes i am asking for the value :) 2^8 = 256 ,right ?

OpenStudy (yacoub1993):

yes

hartnn (hartnn):

so numerator = 256-1 =255 and you sum is just -8*255 =...

OpenStudy (yacoub1993):

-2040

hartnn (hartnn):

that would be correct sum :)

OpenStudy (yacoub1993):

thnxs @hartnn

hartnn (hartnn):

welcome ^_^